When you were born, Claire, I was 24 years old. Today I'm twice the age you were when I was the age you are. What's my age?
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

When you were born, Claire, I was 24 years old.
m = my present age
then
(m-24) = Claire's age
Today I'm twice the age you were when I was the age you are.
2(m-24) = m
2m - 48 = m
2m - m = 48
m = 48 is my age
:
:
Clair is 48-24 = 24, 24 yrs ago m was 24

Answer by ikleyn(52790)

.
When you were born, Claire, I was 24 years old. Today I'm twice the age you were when I was the age you are. What's my age?
~~~~~~~~~~~~~~~~~~~~~~~


Let m = my age (the age of the person, who tells the story).

Then the Clair age is (m-24).


We should accurately interpret this phrase 

    "Today I'm twice the age you were when I was the age you are."


It means that we should start our equation in this way

    m = 2 * something.


This "something" is  "the age you were when I was the age you are."

If you think a minute, you will get the correct idea that this "something" = (m-24)-24 = m - 48.


Therefore, the equation in whole has the form

    m = 2*(m-48).

Its solution is  m = 2*48 = 96.


ANSWER.  m= 96.  In other words, "my" age is 96 years.

. . . . . . . .

You may check it  ON  YOUR  OWN  that my answer is correct.

You also may check,  on your own,  that the answer given by  @ankor@dixie-net.com is wrong.

---------------

This problem relates to the class of  "twisted"  age problems.

If you want to see other similar solved  "twisted"  age problems,  look into the lesson
    - Really intricate age word problems
in this site.